The LD50 for a substance is the dose that would be lethal for 50% of the population that receive that dose. The LD50 for aspirin in rats is 1.75 grams per kilogram of body weight. How many aspirin tablets, each of which contains 325 mg of aspirin, would you have to feed a rat weighing 290. g to give him this dose?
1
Expert's answer
2017-10-03T15:00:07-0400
To reach the LD50 dose, one must give (LD50 times weight of the rat) grams of aspirin:

m=1.75 (g·kg^(-1))·290 (g)=0.5075 (g)
Then, we can calculate the number of aspirin tablets that contain 0.5075 g of aspirin:

N=(0.5075(g))/(325 (mg))=1.56 ≈1.6

Answer: 1.6 tablets